Find the slope and y intercept of each equation. Then graph.
y=x+2
y=4/5x+2
y=7/4x+2

Answer by rapaljer(4551)

When the equation is in slope-intercept form, y=mx+b, the slope is always the coefficient of x, and the y-intercept is always the number term.

y=x+2
Slope = 1, y-intercept = 2.

y=4/5x+2
Slope = 4/5, y-int = 2

y=7/4x+2
Slope = 7/4, y-int = 2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus